OpenStudy (anonymous):

square root12 x square root2=? perform the indicated operation. simplify the answer when possible.

OpenStudy (anonymous):

\[\sqrt{12}\times \sqrt{2}\] \[2\sqrt{3}\times \sqrt{2}\] =\[2\sqrt{3\times2}\] =\[2\sqrt{6}\]